OpenStudy (anonymous):

Find the horizontal limit(s) of the following function: f(x) = \frac {11 x^3 - 9 x^2 -10 x }{ 9 - 11 x - 10 x^3 } ?and ?

OpenStudy (anonymous):

(-11x^3+9x^2+10x)

OpenStudy (anonymous):

how so i solve it?

OpenStudy (shayaan_mustafa):

kindly make standard form of your question. so I could understand.

OpenStudy (shayaan_mustafa):

@rukh

OpenStudy (anonymous):

Find the horizontal limit(s) of the following function: f(x) = \frac {11 x^3 - 9 x^2 -10 x }{ 9 - 11 x - 10 x^3 }

OpenStudy (shayaan_mustafa):

I told in fraction form correctly..

OpenStudy (shayaan_mustafa):

\[f(x)=(11x ^{3}-9x ^{2}-10x)/(9-11x-10x ^{3})\]is this your fraction looks like?

OpenStudy (anonymous):

see attached

OpenStudy (anonymous):

yes

OpenStudy (shayaan_mustafa):

ok now.. as in your question. horizontal limits mean domain of the function. DOMAIN: it is where function is defined. for example if\[\lim_{x \rightarrow 0}1/x=\infty\] because anything divided by zero, goes to infinity. so the function is not defined at infinity. so every number except this is its domain. because at 0 1/x is not defined.

OpenStudy (shayaan_mustafa):

got it ?

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

but the answer will be a set then?

OpenStudy (shayaan_mustafa):

yes. it will be a set.

OpenStudy (anonymous):

horizontal asymptote is the ratio of the leading coefficents, since the degree of the numerator and denominator are the same (they are both 3)

OpenStudy (shayaan_mustafa):

\[f(x)=(11x ^{3}-9x ^{2}-10x)/(9-11x-10^{3})\]In this function denominator must not be zero.

OpenStudy (anonymous):

therefore your horizontal asymptote is \[y=-\frac{11}{10}\]

OpenStudy (anonymous):

ok, my homework program does not accept a set as the answer, it says that the answers must be a number

OpenStudy (anonymous):

and secondly - 11/10 is not being accepted either

OpenStudy (anonymous):

then there is a mistake, but i can assure you that is what it is unless they want you to write \[y=-\frac{11}{10}\]

OpenStudy (anonymous):

im supposed to get 2 answers for this problem

OpenStudy (anonymous):

if -11/10 is one..whats the other?

OpenStudy (mertsj):

Did you double check to see if the problem is posted correctly?

OpenStudy (mertsj):

Or maybe they want it in decimal form.

OpenStudy (anonymous):

yes its correct

OpenStudy (mertsj):

What's correct?

OpenStudy (shayaan_mustafa):

it will define for all real numbers. X=all

OpenStudy (anonymous):

the problem was posted correctly

OpenStudy (anonymous):

i need 2 answers . if -11/ 10 is one, whats the other?

OpenStudy (mertsj):

I don't know. Are they asking for the horizontal asymptotes? Because there is only one.

OpenStudy (anonymous):

there is one, it is the ratio of the leading coefficients. there is no other

OpenStudy (anonymous):

this is so complicated....:(

OpenStudy (mertsj):

Maybe they want you to say as x approaches positive infinity, the limit is -11/10 and as x approaches negative infinity, the limit is -11/10

OpenStudy (anonymous):

they are asking for horizonal limits...is this the same as horizontal asymptotes?

OpenStudy (mertsj):

Satellite???

OpenStudy (anonymous):

ok both answers are -11/10 and they did want it in decimal form. thanks guys

OpenStudy (mertsj):

yw. I'll take the bows for Satellite's work.

OpenStudy (shayaan_mustafa):

as I have solved this equation vertical asymptote= NO horizontal asymptote= -11/10 oblique asymptote= NO

OpenStudy (anonymous):

\[f(x) = \frac {11 x^3 - 9 x^2 -10 x }{ 9 - 11 x - 10 x^3 }\] is what is written, i cannot read the word document. numerator is a polynomial of degree 3 denominator is a polynomial of degree 3 (same degree) horizontal asymptote, limit as x goes to infinity, etc is \[y=-\frac{11}{10}=-1.1\] the ratio of the leading coefficeints. there is no other answer

OpenStudy (anonymous):

thanks so much

OpenStudy (shayaan_mustafa):

you are ever welcome.